OpenStudy (anonymous):

I'm trying to evaluate this limit using L'Hopital's rule... But apparently I'm doing something wrong! Would appreciate if someone could spot what it is I'm doing wrong!

OpenStudy (anonymous):

\[\lim _{x \rightarrow \infty} 6xtan(5/x)\]

OpenStudy (anonymous):

So I rewrote it as\[\lim _{x \rightarrow \infty} \frac{ \tan(5/x) }{ \frac{ 1 }{ 6x } }\]

OpenStudy (anonymous):

Once applying L'Hopital's rule I get... \[\frac{ (\frac{ 5 }{ x^2 })\sec^2(5/x) }{ -36x }\]

OpenStudy (anonymous):

And I believe that the limit of the above equation would be equal to 0... Why is that wrong?

OpenStudy (solomonzelman):

can you use the numerical approach. I would do so here.... I am not coming up with a way to re-write this so that you can differentiate top and bottom.... with numerical approach it seems to approach 30 (I plugged in a couple of values). -------------------------------------------------------

OpenStudy (anonymous):

Hmmm... so just to plug in large numbers?

OpenStudy (solomonzelman):

well, I did it for the check. Just that we know the correct answer. I plugged very large numbers to see how the function behaves the more x tends towards infinity.

OpenStudy (solomonzelman):

I am trying to think f a good way to this one, if there is.

OpenStudy (anonymous):

Mhmm right

OpenStudy (anonymous):

not to butt in but i think your derivative is wrong

OpenStudy (anonymous):

\[\frac{d}{dx}[\frac{1}{6x}]=-\frac{1}{6x^2}\]

OpenStudy (anonymous):

also the numerator is off by a minus sign

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~\infty }~\frac{(-5/x^2)\sec^2\left(\frac{5}{x}\right)}{-\frac{1}{6x^2}}}\) \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~\infty }~\frac{(5/x^2)\sec^2\left(\frac{5}{x}\right)}{\frac{1}{6x^2}}}\) \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~\infty }~30\sec^2\left(\frac{5}{x}\right)}\) yes, should be this.

OpenStudy (anonymous):

that looks better!

OpenStudy (solomonzelman):

in my third line, dividing by 1/6 is multiplying times 6. and 1/x^2 cancel

OpenStudy (solomonzelman):

yes, now plug in that...

OpenStudy (solomonzelman):

30 times (1)^2

OpenStudy (anonymous):

How does the numerator become negative?

OpenStudy (solomonzelman):

because you are taking the derivative of 5/x and that is a negative exponent

OpenStudy (anonymous):

Oh oh oh!!! Yeah just a confusion sorry!

OpenStudy (solomonzelman):

derivat. of (5/x) is the chain rule for your angle....do you get why it is 30 ?

OpenStudy (anonymous):

Im trying to figure that out...

OpenStudy (solomonzelman):

did you get everything up to and including \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~\infty }~30\sec^2\left(\frac{5}{x}\right)}\) ?

OpenStudy (anonymous):

Still thinking about the last step

OpenStudy (solomonzelman):

ok take your time

OpenStudy (anonymous):

Ahhh yeah i got it!!

OpenStudy (solomonzelman):

yes, so then we have: \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~\infty }~30\sec^2\left(\frac{5}{x}\right)}\) applying our limit properties, \(\large\color{slate}{\displaystyle 30\sec^2\left(\lim_{x \rightarrow ~\infty }\frac{5}{x}\right)}\)

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle 30\sec^2\left(0\right)}\)

OpenStudy (anonymous):

Thanks so much for the help @SolomonZelman and @satellite73 :)

OpenStudy (solomonzelman):

yes, and then you know sec^2(0)=1/cos^2(0)=1/1^2=1

OpenStudy (solomonzelman):

so then 30 * 1 ....

OpenStudy (anonymous):

Yes, I understand the rest :) Thanks so much!

OpenStudy (solomonzelman):

yw